# Implicit Bias in Matrix Factorization and its Explicit Realization in a New Architecture

### v1

@@ 4.1.3 Reducing Network Size with SVD-based Pruning
Efficient and lightweight feed-forward layers are crucial for real-world applications. For instance, the Apple Intelligence Foundation Models (Gunter et al., 2024) recently reported that pruning hidden dimensions in feed-forward layers yields the most significant gains in their foundation models. Building on this insight, we leverage the inherent low-rank bias of the UDV architecture through an SVD-based pruning strategy to produce compact networks without sacrificing performance.
@@
$UD=\texttt{USV}^{\top},\quad\texttt{U}\in\mathbb{R}^{d\times m},\quad\texttt{S}\in\mathbb{R}^{m\times m},\quad\texttt{V}^{\top}\in\mathbb{R}^{m\times m}.$ (8)
By dropping small singular values in S, these matrices can be truncated to $\smash{\bar{\texttt{U}}\in\mathbb{R}^{d\times r}}$, $\smash{\bar{\texttt{S}}\in\mathbb{R}^{r\times r}}$ and $\bar{\texttt{V}}^{\top}\in\mathbb{R}^{r\times m}$, where $0<r<m$. Consequently, ($m-r$) neurons can be pruned, and new weight matrices are assigned:
$\bar{U}=\bar{\texttt{U}}\in\mathbb{R}^{d\times r},\quad\bar{D}=\bar{\texttt{S}}\in\mathbb{R}^{r\times r},\quad\bar{V}=\bar{\texttt{V}}^{T}V\in\mathbb{R}^{r\times c}.$ (9)
We applied this pruning strategy on the models from Table 1. The left part of Figure 4 presents an example comparing the generalization capability of pruned models. For comparison, we created compact models by training from scratch with a reduced number of neurons $m$ in the hidden layer. The performance change for these models is shown in the right panel of Figure 4. Although our pruned networks derived from UDV demonstrate that models with significantly fewer parameters can still achieve strong generalization, these compressed architectures are often more challenging to optimize directly within the reduced space, consistent with prior findings in the literature (Arora et al., 2018; Chang et al., 2021). We omit the results for retraining with the UV model, as they show similar trends to UDV in this context, though UDV generally exhibits superior generalization.
Figure 4: Comparison of SVD-based pruning and re-training compact networks on the HPART dataset using the NAdam algorithm with a learning rate of $10^{-3}$. Negative percentages indicate improvements over the baseline. SVD-based pruning demonstrates that the UDV leads to a compact model without performance degradation, while retraining shows that the UDV achieves better generalization in a compact model.
@@ 4.1.4 Further Details and Discussions
Our findings in experiments on neural networks align with the results observed in matrix factorization. A key distinction, however, was the use of different optimization algorithms, including stochastic gradients and momentum steps, in neural network experiments. Despite these differences, the UDV architecture consistently demonstrated a strong bias toward low-rank solutions. Additional experiments and details can be found in the Appendices, with key results summarized below:
  • In the early stages of this research, we explored four variants of UDV, each differing slightly in their constraints. We selected the version presented in (7), as it generally exhibits the most pronounced decay in the singular value spectrum. For completeness, details of the other three variants are provided in Section B.1.
  • Section B.2 provides additional results for the experiment described in Section 4.1.2. Specifically, we present results analogous to those in Table 1 and Figure 3, but focusing exclusively on the MBGDM algorithm. These results exhibit similar trends, reinforcing consistency across different methods. Additionally, comprehensive performance comparisons across all algorithms and models are provided in Tables SM2, SM3, SM4 and SM5 in the Appendices.
  • Additional results on SVD-based pruning are provided in Section B.3. We demonstrate that the UDV framework consistently achieves low-rank solutions across various problem configurations. Furthermore, we analyze the effect of learning rate on the singular value spectrum, similar to the analysis in Figure 1, but applied to neural network experiments. This analysis confirms that the UDV framework produces low-rank solutions across a broad range of learning rates.
  • Section B.4 extends the UDV framework by incorporating ReLU activation. Preliminary experiments with the UDV-ReLU model indicate that it also tends to yield low-rank solutions, similar to those observed in the original UDV framework.
  • Prior work on implicit bias in neural networks suggests that increasing depth enhances the tendency toward low-rank solutions (Arora et al., 2019; Feng et al., 2022), raising the question of whether the pronounced bias in the UDV framework is just a consequence of adding a diagonal layer. To investigate this, we conducted experiments comparing the UDV model to fully connected three-layer networks, as detailed in Section B.5. Additionally, we included a UDV model without constraints in these comparisons. The results indicate that this bias cannot be attributed solely to depth, highlighting the critical role of explicit constraints.
  • Section B.6 compares the spectral decay of the UDV network to that induced by classical weight decay regularization in two- and three-layer networks. The results show that while weight decay promotes singular value decay, it can not reproduce the strong decay observed in the UDV model.
@@ 5 Conclusions
We proposed a new matrix factorization framework, inspired by the observation that implicit bias is driven by dynamics—potentially divergent—that are distinct from those leading to convergence of the objective function. This framework constrains the factors within Euclidean norm balls and introduces a middle diagonal factor to ensure the search space is not restricted. Numerical experiments demonstrate that this approach significantly strengthens the low-rank bias in the solution.
To explore the broader applicability of our findings, we designed an analogous neural network architecture with three layers, constraining the fully connected layers and adding a diagonal hidden layer, referred to as UDV. Extensive experiments show that the proposed UDV architecture achieves competitive performance compared to standard fully connected networks, while inducing a structured solution with a strong bias toward low-rank representations. Additionally, we explored the utility of this low-rank structure by applying an SVD-based pruning strategy, illustrating how it can be leveraged to construct compact networks that are more efficient for downstream tasks.
@@ A.1 Matrix Completion under Gaussian Noise
A.2 UDU Factorization Produces Rank-Revealing Solutions
In Figure SM2, we illustrate the evolution of the column norms of $U$ from the matrix completion experiment described in Section 3.2. Figure SM3 displays the corresponding entries in the diagonal factor $D$. For comparison, Figure SM4 shows the column norms of $U$ obtained from the same experiment using the standard BM factorization. The results are shown for $\xi=10^{-1}$ and $\eta=10^{-1}$ over $10^{5}$ iterations, but the findings are similar across other parameter settings. We also repeat the experiment with the noisy measurements described in Section A.1. The results are shown in Figures SM5, SM6 and SM7.
Figure SM2: Evolution of the column norms of $U$ during the matrix completion experiment using the UDU factorization. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
Figure SM3: Evolution of the diagonal entries of $D$ during the matrix completion experiment using the UDU factorization. The x-axis represents indices.
Figure SM4: Evolution of the column norms of $U$ during the matrix completion experiment using the standard BM factorization, shown for comparison. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
Figure SM5: Evolution of the column norms of $U$ during the matrix completion experiment using the UDU factorization with noisy measurements. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
Figure SM6: Evolution of the diagonal entries of $D$ during the matrix completion experiment using the UDU factorization with noisy measurements. The x-axis represents indices.
Figure SM7: Evolution of the column norms of $U$ during the matrix completion experiment using the standard BM factorization with noisy measurements, shown for comparison. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
A.3 Phase Retrieval Image Recovery
We conduct a numerical experiment on the matrix sensing problem arising in phase retrieval image recovery (Candes et al., 2013). The objective is to recover a signal from its quadratic measurements of the form $y_{i}=|\langle a_{i},x\rangle|^{2}$. Although the standard maximum likelihood estimators lead to a non-convex optimization problem due to the quadratic terms, the problem can be reformulated as minimizing a convex function under a rank constraint in a lifted space. By denoting the lifted variable as $X=xx^{\top}$, the measurements can be expressed as
@@ A.3 Phase Retrieval Image Recovery
$\smash{\min_{X\in\mathbb{S}_{+}^{d\times d}}\quad\frac{1}{2}\|\mathcal{A}(X)-b\|_{2}^{2}\quad\mathrm{subj.\ to}\quad\mathrm{rank}(X)\leq 1,}$
representing a special case of problem (1).
@@
In this experiment, we use a gray-scale image of size $16\times 16$ pixels, selected from the Pixel Art dataset (Elgazar, 2024), as the signal $x\in\mathbb{R}^{n}$ to recover, where $n=256$. We generate a synthetic measurement system by sampling $a_{1},\ldots,a_{m}$ from the standard Gaussian distribution, with $m=2n$. We then solve problems (2) and (4). We initialized $U_{0}\in R^{n\times n}$ with entries drawn iid from the standard Gaussian distribution, then rescaled to have a unit Frobenius norm. For $D_{0}\in R^{n\times n}$, we used the identity matrix. We used step-size $\eta=\frac{1}{L}$ where $L$ denotes the smoothness constant. After solving the problem, we recover $x\in\mathbb{R}^{n}$ from the lifted variable $X\in\mathbb{R}^{n\times n}$ by selecting its dominant eigenvector.
@@
Figure SM8 demonstrates that the proposed method consistently identifies low-rank solutions, in line with the results of our other experiments. Figure SM9 displays the recovered image, demonstrating that the proposed method achieves higher quality than BM factorization within the same number of iterations, which can be attributed to the low-rank structure of the solution.
@@
Figure SM8: Comparison of singular value spectrum in recovered image between methods based on $UDU^{T}$ and $UU^{T}$.
Figure SM9: Comparison of recovered image between methods based on $UDU^{T}$ and $UU^{T}$.
@@ Appendix B Additional Details on Neural Network Experiments
B.1 Design Variants of the UDV Architecture

### v2

@@ 3.1.3 Reducing Network Size via SVD-based Pruning
Efficient and lightweight feed-forward layers are crucial for real-world applications. For instance, the Apple Intelligence Foundation Models (Gunter et al., 2024) recently reported that pruning hidden dimensions in feed-forward layers yields the most significant gains in their foundation models. Building on this insight, we leverage the low-rank bias of the UDV architecture through an SVD-based pruning strategy to produce compact networks without sacrificing performance.
@@
$UD=\texttt{USV}^{\top},\quad\texttt{U}\in\mathbb{R}^{d\times m},\quad\texttt{S}\in\mathbb{R}^{m\times m},\quad\texttt{V}^{\top}\in\mathbb{R}^{m\times m}.$
By dropping small singular values in S, these matrices can be truncated to $\smash{\bar{\texttt{U}}\in\mathbb{R}^{d\times r}}$, $\smash{\bar{\texttt{S}}\in\mathbb{R}^{r\times r}}$ and $\bar{\texttt{V}}^{\top}\in\mathbb{R}^{r\times m}$, where $0<r<m$. Consequently, (mrm-r) neurons can be pruned, and new weight matrices are assigned:
$\bar{U}=\bar{\texttt{U}}\in\mathbb{R}^{d\times r},\quad\bar{D}=\bar{\texttt{S}}\in\mathbb{R}^{r\times r},\quad\bar{V}=\bar{\texttt{V}}^{\top}V\in\mathbb{R}^{r\times c}.$
The bottom row of Figure 5 shows test accuracy of the post-training pruned models as a function of the number of remaining neurons. The x-axis is cropped for clarity (the full model has 2520 neurons before pruning). The UDV-based models maintain high accuracy even under aggressive pruning and consistently outperform the UV baseline at comparable compression levels. This demonstrates that UDV induces low-rank representations that are more amenable to compression without compromising performance. Although the pruned networks show that models with significantly fewer parameters can still generalize well, training such small models from scratch typically results in lower test accuracy, consistent with prior findings in the literature (Arora et al., 2018; Chang et al., 2021).
Figure 5: Singular value spectra and post-training SVD-based pruning results for RegNetX-32GF on CIFAR-100 under different optimization algorithms. [Top] UDV induces faster spectral decay than the baseline UV model. [Bottom] Test accuracy as a function of remaining neurons after pruning (full model has 2520 neurons, x-axis cropped for clarity). UDV-based networks retain high accuracy under aggressive pruning
@@ 3.1.4 Further Details and Discussions
Our findings in experiments on neural networks align with the results observed in matrix factorization. A key distinction, however, was the use of different optimization algorithms, including stochastic gradients and momentum steps, in neural network experiments. Despite these differences, the UDV architecture consistently demonstrated a strong bias toward low-rank solutions. Additional experiments can be found in the Appendices, with key results summarized below.
Section B.1 provides comprehensive results complementing those in Sections 3.1.2 and 3.1.3, covering other configurations (MaxViT-T and EfficientNet-B0) and datasets (CIFAR-10/100). The low-rank patterns persist across settings, and the performance gap in pruning between UDV and UV becomes more pronounced as task complexity increases, highlighting the growing advantage of low-rank structure.
Early in this work we explored four UDV variants with different constraints. All produced low-rank solutions, but the formulation in (5) showed the strongest spectral decay, and we adopt it as the main model. The remaining variants are described in Section B.2. Additional experiments on SVD-based pruning with these UDV variants are provided in Section B.3. Furthermore, we analyzed the effect of learning rate on the singular value spectrum, similar to the analysis in Figure 1, but applied to neural network experiments. This analysis confirms that the UDV framework produces low-rank solutions across a broad range of learning rates.
Section B.4 extends the UDV framework by incorporating ReLU activation. Preliminary experiments suggest that UDV-ReLU also promotes low-rank solutions, consistent with the original UDV model.
Prior work on implicit bias in neural networks suggests that network depth can enhances the tendency toward low-rank solutions (Arora et al., 2019; Feng et al., 2022), raising the question of whether the pronounced bias in UDV is just a consequence of the extra diagonal layer. To test this, we compared UDV with fully connected three-layer networks in Section B.5. We also examined an unconstrained UDV variant. The results show that the bias cannot be explained by depth alone, highlighting the importance of explicit constraints.
Section B.6 compares the spectral decay of the UDV network to that induced by classical weight decay regularization in two- and three-layer networks. While weight decay promotes singular value decay, it can not reproduce the strong decay observed in the UDV model.
Finally, Section B.7 presents a toy example applying the UDV structure within the LoRA framework to fine-tune a pre-trained LLaMA-2 model on a causal language modeling task. As a proof of concept, the pruning results suggest that UDV blocks could serve as drop-in replacements for linear layers, showing promising performance under compression.
Most importantly, our comprehensive set of experiments shows that the UDV framework consistently induces a strong implicit bias toward low-rank solutions across diverse architectures, datasets, training setups and algorithms.
@@ 4 Conclusions
We introduced a new matrix factorization framework that constrains the factors within Euclidean norm balls and incorporates a diagonal middle factor to avoid restricting the search space. Numerical experiments show that this formulation enhances the implicit bias toward low-rank solutions.
To explore the broader applicability of our findings, we designed an analogous neural network architecture with three layers, constraining the fully connected layers and adding a diagonal hidden layer, referred to as UDV. Extensive experiments (including those in Appendices) show that the UDV architecture induces a strong bias toward low-rank representations. Additionally, we demonstrated the utility of this low-rank structure by applying an SVD-based pruning strategy, as a proof of concept, illustrating how it can be leveraged to construct compact networks that are more efficient for downstream tasks.
The proposed model exhibits reduced rank regression behavior, where the training process gradually increases the model rank. We believe these our should be of broader interest to research on implicit bias. Although we provide some theoretical insights, developing a more complete theory and designing algorithms that fully exploit the model’s regularization capabilities remain important directions for future work.
@@ A.1 Matrix Completion under Gaussian Noise
A.2 Phase Retrieval
A.2.1 Phase Retrieval on Synthetic Data
The experiment with real data in Section 2.2.2 is large-scale and thus impractical for testing full-dimensional factorizations with $r=d$. To complement it, we perform a small-scale synthetic test bed for phase retrieval.
We consider the matrix sensing formulation of phase retrieval (Candes et al., 2013), where the goal is to recover a signal from quadratic measurements of the form $y_{i}=|\langle a_{i},x\rangle|^{2}$. Although the standard maximum likelihood estimators lead to a non-convex optimization problem due to the quadratic terms, the problem can be reformulated as minimizing a convex function under a rank constraint in a lifted space. By denoting the lifted variable as $X=xx^{\top}$, the measurements can be expressed as
@@ A.2.1 Phase Retrieval on Synthetic Data
$\smash{\min_{X\in\mathbb{S}_{+}^{d\times d}}\quad\frac{1}{2}\|\mathcal{A}(X)-b\|_{2}^{2}\quad\mathrm{s.t.}\quad\mathrm{rank}(X)\leq 1.}$
@@
In this experiment, we use a gray-scale image of size $16\times 16$ pixels, selected from the Pixel Art dataset (Elgazar, 2024), as the signal $x\in\mathbb{R}^{n}$ to recover, where $n=256$. We generate a synthetic measurement system by sampling $a_{1},\ldots,a_{m}$ from the standard Gaussian distribution, with $m=2n$. We then solve problems (1) and (2). We initialized $U_{0}\in\mathbb{R}^{n\times n}$ with entries drawn iid from the standard Gaussian distribution, then rescaled to have a unit Frobenius norm. For $D_{0}\in\mathbb{R}^{n\times n}$, we used the identity matrix. We used step-size $\eta=\frac{1}{L}$ where $L$ denotes the smoothness constant. After solving the problem, we recover $x\in\mathbb{R}^{n}$ from the lifted variable $X\in\mathbb{R}^{n\times n}$ by selecting its dominant eigenvector.
@@
Figure SM2 demonstrates that the proposed method consistently identifies low-rank solutions, in line with the results of our other experiments. Figure SM3 displays the recovered image, demonstrating that the proposed method achieves higher quality than BM factorization within the same number of iterations, which can be attributed to the low-rank structure of the solution.
@@
Figure SM2: Comparison of singular value spectrum in recovered image between methods based on $UDU^{T}$ and $UU^{T}$.
Figure SM3: Comparison of recovered image between methods based on $UDU^{T}$ and $UU^{T}$.
A.2.2 Intermediate Reconstructions in Fourier Ptychography Imaging
We provide additional details for the Fourier ptychography imaging experiment in Section 2.2.2, showing the reconstructions at intermediate iterations (Figure SM4).
$k=0$
$k=10$
$k=100$
$k=1000$
$k=10000$
Figure SM4: Reconstructed images obtained from the proposed UDU factorization (top row) and the standard BM factorization (bottom row) at different iterations, starting from the same random initialization. While the UDU method converges to a clean and accurate recovery, the BM reconstructions display severe artifacts.
A.3 UDU Factorization Produces Rank-Revealing Solutions
In Figure SM5, we illustrate the evolution of the column norms of $U$ from the matrix completion experiment described in Section 2.2. Figure SM6 displays the corresponding entries in the diagonal factor $D$. For comparison, Figure SM7 shows the column norms of $U$ obtained from the same experiment using the standard BM factorization. The results are shown for $\xi=10^{-1}$ and $\eta=10^{-1}$ over $10^{5}$ iterations, but the findings are similar across other parameter settings. We also repeat the experiment with the noisy measurements described in Section A.1. The results are shown in Figures SM8, SM9 and SM10.
Figure SM5: Evolution of the column norms of $U$ during the matrix completion experiment using the UDU factorization. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
Figure SM6: Evolution of the diagonal entries of $D$ during the matrix completion experiment using the UDU factorization. The x-axis represents indices.
Figure SM7: Evolution of the column norms of $U$ during the matrix completion experiment using the standard BM factorization, shown for comparison. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
Figure SM8: Evolution of the column norms of $U$ during the matrix completion experiment using the UDU factorization with noisy measurements. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
Figure SM9: Evolution of the diagonal entries of $D$ during the matrix completion experiment using the UDU factorization with noisy measurements. The x-axis represents indices.
Figure SM10: Evolution of the column norms of $U$ during the matrix completion experiment using the standard BM factorization with noisy measurements, shown for comparison. The x-axis represents column indices, and the y-axis shows the Euclidean norms of the columns.
A.4 Additional Details on the Fixed Point Analysis
In Section 2.3, we provided insights into the inner workings of the UDU framework on matrix factorization problems based on a fixed-point analysis. Recall the algorithm steps:
$U_{k+1}$ $=\Pi_{U}(\bar{U}_{k+1})$ $\text{where}\quad\bar{U}_{k+1}=U_{k}-2\eta\nabla f(U_{k}D_{k}U_{k}^{\top})U_{k}D_{k}$
$D_{k+1}$ $=\Pi_{D}(\bar{D}_{k+1})$ $\text{where}\quad\bar{D}_{k+1}=D_{k}-\eta U_{k}^{\top}\nabla f(U_{k}D_{k}U_{k}^{\top})U_{k}$
Let $(U,D)$ denote a fixed point of this algorithm and let $X=UDU^{\top}$. Suppose $\lVert\bar{U}\rVert\leq\alpha$. In this case, the projection step $\Pi_{U}$ acts as an identity:
$U=\bar{U}=U-2\eta\nabla f(X)UD\quad\implies\quad\nabla f(X)UD=0$ (6)
Otherwise, if we assume $\lVert\bar{U}\rVert\geq\alpha$, then the projection $\Pi_{U}$ has a normalization effect:
$U=\frac{\alpha}{\lVert\bar{U}\rVert}\bar{U}=\frac{\alpha}{\lVert\bar{U}\rVert}\big(U-2\eta\nabla f(X)UD\big)$
which we can reorganize as
$\nabla f(X)UD=-\beta U\quad\text{where}\quad\beta:=\frac{\lVert\bar{U}\rVert-\alpha}{2\alpha\eta}.$ (7)
Denoting the columns of $U$ by $u_{j}$ and the diagonal entries of $D$ by $\lambda_{j}$, we express the conditions (6) and (7) in terms of the individual components as presented in conditions $(a)$ and $(b)$ in Section 2.3.
Next, we investigate the $D$ component. Note that the projection $\Pi_{D}$ operates in an entry-wise separable manner; it maps all non-diagonal entries to 0, and takes the positive part for the diagonal entries. It maps all off-diagonal entries to zero and applies the positive-part operator to the diagonal entries. Hence, it suffices to examine only the diagonal elements.
Suppose $\bar{\lambda}_{j}\leq 0$. Then $\lambda_{j}=\max(\bar{\lambda}_{j},0)=0$, and
$\bar{\lambda}_{j}=\lambda_{j}-\eta\,u_{j}^{\top}\nabla f(X)u_{j}=-\eta\,u_{j}^{\top}\nabla f(X)u_{j}\quad\implies\quad u_{j}^{\top}\nabla f(X)u_{j}\geq 0.$ (8)
Otherwise, if $\bar{\lambda}_{j}>0$, then $\lambda_{j}=\max(\bar{\lambda}_{j},0)=\bar{\lambda}_{j}$, hence
$\lambda_{j}=\bar{\lambda}_{j}=\lambda_{j}-\eta\,u_{j}^{\top}\nabla f(X)u_{j}\quad\implies\quad u_{j}^{\top}\nabla f(X)u_{j}=0.$ (9)
The conditions in (8) and (9) correspond to cases $(c)$ and $(d)$ in Section 2.3.
@@ Appendix B Additional Details on Neural Network Experiments
Computing environment. All classification tasks were conducted on an NVIDIA A100 GPU with four cores of the AMD Epyc 7742 processor, while regression tasks were conducted on a single core of an Intel Xeon Gold 6132 processor. We used Python 3.12.3 and PyTorch 2.5.1.
Robustness Evaluation. All reported results in the tables were averaged across 1000 seeds on HPART, 100 on NYCTTD, and 10 on CIFAR-10, CIFAR-100, and MNIST.
B.1 Full Results of Neural Networks Reported in the Main Text
Here we provide all results related to the image classification experiments discussed in Section 3. Table SM1 and Table SM2 present the complete results corresponding to Table 1 in the main text, while Figures SM11, SM12, SM13, SM14, SM15 and SM16 show all singular value spectrum and pruning results. The regression experiments were part of the earlier stage of this study, and their details are reported in Section B.2. Meanwhile, the early experiments reduced the need for extensive hyperparameter selection in the current study. Therefore, for Adam and NAdam we used learning rates of $10^{-3}$ and $10^{-4}$, whereas for MBGD and MBGDM we used $10^{-1}$ and $10^{-2}$.
Table SM1: Comparison of model performance on CIFAR-10. Note “–” denotes results that are not applicable.
Model MaxVit-T [M] EfficientNet-B0 [E] RegNetX-32GF [R]
Optimizer Adam NAdam MBGD MBGDM Adam NAdam MBGD MBGDM Adam NAdam MBGD MBGDM
LR: $10^{-4}$
UDV:99.14
UV:99.11
UDV:99.17
UV:99.11
UDV:97.87
UV:97.98
UDV:97.86
UV:97.99
UDV:98.67
UV:98.64
UDV:98.55
UV:98.56
LR: $10^{-3}$
UDV:98.61
UV:98.76
UDV:98.72
UV:98.69
UDV:97.61
UV:97.77
UDV:97.68
UV:97.69
UDV:97.84
UV:97.60
UDV:97.78
UV:97.68
LR: $10^{-2}$ - -
UDV:98.04
UV:98.25
UDV:98.93
UV:98.94
UDV:95.98
UV:96.65
UDV:97.79
UV:97.90
UDV:98.00
UV:98.18
UDV:98.61
UV:98.65
LR: $10^{-1}$ - -
UDV:98.93
UV:99.01
UDV:99.05
UV:99.03
UDV:97.86
UV:97.82
UDV:97.87
UV:97.94
UDV:98.67
UV:98.64
UDV:98.59
UV:98.63
Table SM2: Comparison of model performance on CIFAR-100. Note “–” denotes results that are not applicable and $\times$” denotes results that did not converge.
Model MaxVit-T [M] EfficientNet-B0 [E] RegNetX-32GF [R]
Optimizer Adam NAdam MBGD MBGDM Adam NAdam MBGD MBGDM Adam NAdam MBGD MBGDM
LR: $10^{-4}$
UDV:84.69
UV:87.75
UDV:84.69
UV:87.72
UDV:91.58
UV:91.90
UDV:91.63
UV:92.03
UDV:89.17
UV:89.22
UDV:89.31
UV:89.40
LR: $10^{-3}$
UDV:84.70
UV:86.36
UDV:84.76
UV:86.09
UDV:89.05
UV:90.55
UDV:89.49
UV:90.60
UDV:85.41
UV:85.62
UDV:85.73
UV:85.44
LR: $10^{-2}$ - -
UDV: $\times$
UV:84.57
UDV:85.23
UV:87.73
UDV: $\times$
UV:88.23
UDV:89.20
UV:91.55
UDV: $\times$
UV:88.82
UDV:89.75
UV:90.33
LR: $10^{-1}$ - -
UDV:85.45
UV:87.74
UDV:86.18
UV:86.67
UDV:89.88
UV:91.56
UDV:91.35
UV:91.70
UDV:89.79
UV:90.30
UDV:88.53
UV:88.69
Figure SM11: Singular value spectrum and pruning performance of MaxVit-T on CIFAR-10.
Figure SM12: Singular value spectrum and pruning performance of EfficientNet-B0 on CIFAR-10.
Figure SM13: Singular value spectrum and pruning performance of RegNetX-32GF on CIFAR-10.
Figure SM14: Singular value spectrum and pruning performance of MaxVit-T on CIFAR-100. The learning rate of $10^{-2}$ for MBGD was unsuitable for our model, but no other hyperparameters were fine-tuned to ensure fairness.
Figure SM15: Singular value spectrum and pruning performance of EfficientNet-B0 on CIFAR-100. The learning rate of $10^{-2}$ for MBGD was unsuitable for our model, but no other hyperparameters were fine-tuned to ensure fairness.
Figure SM16: Singular value spectrum and pruning performance of RegNetX-32GF on CIFAR-100. The learning rate of $10^{-2}$ for MBGD was unsuitable for our model, but no other hyperparameters were fine-tuned to ensure fairness.
B.2 Design Variants of the UDV Architecture